A sample of water was alkaline to both phenolphthalein and methyl orange.100 ml of this water sample required 10ml of N/50 h2so4 for phenolphthalein end point.determine the type and extent of alkalinity present


1
Expert's answer
2021-01-20T03:18:16-0500

H2SO4 is dibasic acid so N\50 = 0.02N

Molarity of acid = 0.02\2 = 0.01M

So mol of acid = 10mL × 0.01M

= 0.1mmol

So amount of base = 2×0.1 mmol = 0.2mmol

Conc of base = (0.2×10-3 mol × 1000)\100

=2×10-3 M

POH = -log[OH-] = -log (2×10-3) =2.7

PH + POH = 14

PH = 14 -POH = 14 -2.7 = 11.3 (ANS)